# Conceptual Framework Toward Embodied Collective Adaptive Intelligence

### v1

@@ Abstract
Collective Adaptive Intelligence (CAI) represent a transformative approach in artificial intelligence, wherein numerous autonomous agents collaborate, adapt, and self-organize to navigate complex, dynamic environments. This paradigm is particularly impactful in embodied AI applications, where adaptability and resilience are paramount Fan and Liu (2025). By enabling systems to reconfigure themselves in response to unforeseen challenges, CAI facilitate robust performance in real-world scenarios. This article introduces a conceptual framework for designing and analyzing CAI. It delineates key attributes including task generalization, resilience, scalability, and self-assembly, aiming to bridge theoretical foundations with practical methodologies for engineering adaptive, emergent intelligence. By providing a structured foundation for understanding and implementing CAI, this work seeks to guide researchers and practitioners in developing more resilient, scalable, and adaptable AI systems across various domains.
@@ 1 Prospectives and Challenges
Emergence and Collectivity in Complex Systems: The seminal concept of "More is Different" Anderson (1972) posits that collective behavior underpins a spectrum of phenomena spanning physics, biology, and artificial intelligence Benjamin et al. (2023). This principle—wherein complex patterns emerge from the coordinated interactions of simpler constituents—has profoundly influenced modern deep learning research Ha and Tang (2022). By abstracting collective dynamics, recent advances have sought to reconcile microscopic rules with macroscopic functionalities, such as in neural modular architectures or swarm intelligence models.
Insights from Occam’s Razor: Occam’s Razor, which prioritizes simplicity in explanatory frameworks, offers a complementary perspective on collective systems. In machine learning, models adhering to this principle often demonstrate superior generalization by mitigating overfitting to training noise Domingos (1999), Mirshekari (2023). When applied to collectivity, this heuristic suggests that systems composed of repetitive, parameter-efficient modules may outperform monolithic architectures with comparable or greater parameter counts. Such designs inherently balance expressivity and inductive bias, enabling robust performance across diverse tasks while avoiding the pitfalls of excessive algorithmic complexity.
Computational Challenges in Collective Intelligence: A critical bottleneck in scaling collective intelligent system lies in their architectural divergence from mainstream computational paradigms. While collective models may reduce absolute parameter counts, their computational demands often escalate due to recurrent interactions among components Ha and Tang (2022). This discrepancy arises from parameter interdependencies and sequential binding operations, which limit parallelization on hardware accelerators like GPUs. Consequently, optimizing collective systems necessitates rethinking not only algorithmic design but also co-adapting computational substrates to accommodate their unique operational patterns.
Collective Adaptivity: Beyond Training-Time Constraints: The concept of collective adaptivity—a paradigm emphasizing decentralized, self-organizing, self-adapting multi-agent systems—has garnered significant attention in societal systems Galesic et al. (2023) and software engineering Bucchiarone (2019), Bucchiarone and Mongiello (2019). However, its application within artificial intelligence (AI) remains relatively underexplored. This work adopts relevant definitions and posits that collective adaptive intelligence is capable of dynamically adjusting its size, topology, and functionality during inference—far surpassing the scale anticipated at training or design time. Leveraging collective adaptivity could fundamentally reshape AI architectures by decentralizing learning processes and enabling emergent, context-aware behaviors. Preliminary evidence for this capability emerges from transformative domains such as large language models, which demonstrate emergent collaboration abilities even in absense of such training Hong et al. (2024), and autonomous agents operating in open-ended environments Bauer et al. (2023). However, as of now, the emergence of collective adaptivity remains limited. Emergent behaviors that are markedly distinct from individual behaviors have yet to manifest in practice. Moreover, achieving true collective scalability requires addressing both algorithmic robustness (e.g., graceful degradation under component failure) and architectural flexibility (e.g., dynamic resource allocation), thereby bridging the gap between theoretical emergence and practical deployment.
This paper formally introduces a definition of embodied collective adaptive intelligence (CAI), delineating their defining attributes and outlining a systematic conceptual framework to contextualize their design and analysis. By establishing this foundational structure, we aim to catalyze future advancements in artificial general intelligence (AGI) research, bridging theoretical insights with scalable, interdisciplinary methodologies for engineering adaptive, emergent intelligence.
@@ 2 Collective Adaptive Intelligence: Conceptual Definitions
We first consider a conceptual definition of collective adaptive intelligent (CAI) system. Canonical Multi-Agent Systems (MAS) are trained on single tasks and deployed solely on the training task Weiss (1999). In traditional settings, the coordination patterns among components are predefined, which is regarded as specialist collective intelligence (SCI). However, if the MAS is pretrained on a diverse set of numerous tasks, it can potentially be deployed not only on a broad range of in-distribution tasks but also on some out-of-distribution (OOD)tasks. In such cases,it is considered as a generalist collective intelligence (GCI). Drawing inspiration from some related studies Fernandez-Marquez et al. (2013), Capodieci et al. (2016), CAI should possess at least the following features (as shown in Figure 1):
@@
Adaptation (CAI) vs. Zero-Shot Generalization (GCI): The distinction between adaptation and zero-shot generalization is frequently overlooked in discussions of model generalization. However, in partially observable Markov decision processes (POMDPs), zero-shot generalization is inherently constrained, as established in prior work on meta-learning and general-purpose systems Thrun and Pratt (1998), Kirsch et al. (2023), Fan and Liu (2025). Notably, in collective systems—where individual agents operate with limited environmental visibility—the reliance on inter-agent communication exacerbates this challenge. Here, generalization often demands adaptation to partial observability, which can directly undermine zero-shot performance. This interplay underscores a critical tension: optimizing for zero-shot capabilities may conflict with achieving robust generalization in decentralized, communication-dependent systems. Clarifying this distinction is essential for designing scalable AI frameworks in uncertain or dynamically varying environments.
@@ 3 A Conceptual Framework for Collective Adaptive Agents
We now present a conceptual framework for laying the groundwork of the foundation components for Collective Adaptive Intelligence (CAI), specifically the Collective Adaptive Agents (CAA).
We consider partial Markov decision processes $\tau=\langle\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{T},\mathcal{Z},\mathcal{R},\gamma\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ denotes the action space, $\mathcal{O}$ represents the observation space, $\mathcal{R}$ represents the utility or feedbacks, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ denotes the state transition, $\mathcal{Z}:\mathcal{S}\rightarrow\mathcal{O}$ is the observation model. The target is to maximize the accumulated utility $G(\pi)=\sum_{t}\gamma^{t}r_{t}$ through decision-making $a_{t}\sim\pi$.
We now examine the policy of a CAI system denoted as $\pi_{C}$, which comprises a set of agents $C=\{\mathcal{G}_{1},...,\mathcal{G}_{N}\}$. Each agent $\mathcal{G}_{i}$ is defined as $\langle f,h,\theta,\phi_{i,0},g_{i,0}\rangle$, where $f$ and $h$ represent predefined models and operations, while $\theta$, $\phi_{i,0}$, and $g_{i,0}$ are the parameters subject to optimization. Within this framework, each agent processes a portion of the observation $o_{i,t}$, where $o_{t}\in\mathcal{O}$ and $o_{t}=(o_{1,t},...,o_{n,t})$, and accordingly outputs a corresponding segment of the action $a_{i,t}$, with $a_{t}\in\mathcal{A}$ and $a_{t}=(a_{1,t},...,a_{n,t})$.
We now propose a conceptual framework of collective adaptive agents (CAAs) as following:
$\phi_{i,t},a_{i,t},m_{i,t},g_{i,t}=$ $f(o_{i,t},a_{i,t-1},r_{t-1},M_{i,t-1},\phi_{i,t-1},g_{i,t-1},\theta)$
$M_{i,t}=$$&\displaystyle\{(m_{j,t},g_{j,t})|\forall j,h(g_{i,t},g_{j,t})\leq h_{0}\text{ \& }$
$\text{RANK}_{j}(h(g_{i,t},g_{j,t}))<K\}$ (1)
@@
Figure 2: A formulation of collective adaptive agents.
Equation 1 is characterized by the following features:
  • Decentralization: Equation 1 operates in a decentralized fashion.
  • Self-adapting: By observing $o_{i,t},a_{i,t-1},M_{i,t-1},r_{t-1}$ and updating $\phi_{i,t}$, each individual agent can perform self-adaptation, such as via reinforcement learning techniques, thereby adapting to specific tasks and topologies during testing.
  • Self-assembly and Collective Scalability: By utilizing $h$ and updating $g_{i,t}$, the system can reshape communication and collaboration patterns. Additionally, the number of agents can be adjusted to form systems of varying scales.
The function $f$ can be implemented using recurrent neural structures, Transformers, or linear attention mechanisms. This design allows the agent to function as an in-context learner from the outset. Additionally, $f$ can be interpreted as artificially designed optimization techniques, such as gradient descent with a specified target function. This interpretation transforms Equation 1 into a gradient-based adaptation mechanism. For the function $h$, a suitable choice might be Euclidean distance. In such a scenario, $g_{i,t}$ can be interpreted as the position (estimation) of agent $i$ in a Euclidean space. This setup draws an interesting parallel to the spatial distance between ants in a colony or between neurons in a neural system, highlighting the agent’s position and interaction dynamics within its operational space.
4 Limitations and Potential Challenges
At this stage, we propose a conceptual framework for CAI and its CAA components, though it currently lacks robust validation. Although collective intelligence has strong natural analogies and great potential for the development of artificial intelligence, mainstream computational infrastructure struggles to optimize collective intelligent systems effectively, as noted by Ha and Tang (2022). However, there is promise in exploring whether we can develop genuinely collectively scalable CAAs. The goal is to find systems that can be pre-trained at small scales or even via single-agent training, yet are capable of self-organization and cooperation at much larger scales. While prior advances in LLMs and open-world research offer some encouragement, significant further progress is still required in this area.

### v2

@@ Abstract
Collective Adaptive Intelligence (CAI) represent a transformative approach in embodied AI, wherein numerous autonomous agents collaborate, adapt, and self-organize to navigate complex, dynamic environments. By enabling systems to reconfigure themselves in response to unforeseen challenges, CAI facilitate robust performance in real-world scenarios. This article introduces a conceptual framework for designing and analyzing CAI. It delineates key attributes including task generalization, resilience, scalability, and self-assembly, aiming to bridge theoretical foundations with practical methodologies for engineering adaptive, emergent intelligence. By providing a structured foundation for understanding and implementing CAI, this work seeks to guide researchers and practitioners in developing more resilient, scalable, and adaptable AI systems across various domains.
@@ 1 Prospectives and Challenges
This section explores the conceptual and practical foundations required for building embodied collective adaptive intelligence (CAI). We examine how principles of emergence, modular simplicity, and system-level adaptivity intersect, and what computational and architectural challenges must be overcome to translate these ideas into scalable, real-world AI systems.
Emergence and Collectivity in Complex Systems: The idea that “more is different” [1] has long explained how complex behaviors can arise from the interactions of simple components. This principle—central to phenomena in physics, biology, and artificial intelligence—also underpins recent advances in deep learning. In collective systems such as neural modular architectures and swarm intelligence, researchers have begun to bridge local interaction rules with emergent, system-wide capabilities [7]. These examples show how collectivity can serve as both a functional and scalable design paradigm.
Simplicity as a Design Principle: Occam’s Razor reminds us that simplicity can drive generalization. In machine learning, simpler models often avoid overfitting and adapt better to new inputs [4]. This insight applies to collective systems as well: architectures built from small, reusable, parameter-efficient modules may outperform large monolithic models. By promoting modularity and inductive bias, such systems achieve flexibility without sacrificing robustness—especially important when scaling across diverse tasks and domains.
Why Collective Intelligence Stresses Today’s Hardware: Despite their promise, collective systems face a fundamental mismatch with existing computing infrastructure. Unlike feedforward models that scale linearly on GPUs, collective architectures often rely on recursive message passing and interdependent modules, making them difficult to parallelize [7]. As a result, these models can be more memory- and compute-intensive—even if they use fewer total parameters. Unlocking their potential will require co-designing new hardware and software stacks optimized for decentralized, interaction-heavy computation.
From Collective Intelligence to Collective Adaptivity: Collective adaptivity goes a step further, enabling AI systems to reorganize themselves—changing their structure, behavior, or scale—at inference time. This concept, well-established in complex systems and multi-agent software engineering [6, 3], remains underexplored in AI. We argue that embodied AI systems capable of adapting their topology, communication patterns, or component roles on the fly could break free from the constraints of static training-time architectures.
Early signs of this potential appear in domains like large language models—where emergent task coordination can arise without explicit programming [8]—and in autonomous agents navigating open-ended environments [2]. Still, true collective adaptivity remains rare. Most current systems lack the robustness to degrade gracefully under partial failure, or the flexibility to reallocate resources dynamically in response to context. Bridging this gap between theoretical adaptivity and practical deployment is a major unsolved challenge.
In this article, we formally define embodied collective adaptive intelligence (CAI), and outline its core attributes, such as scalability, decentralization, and self-organization. Our goal is to offer a conceptual framework that supports not just analysis, but the engineering of real-world systems capable of emergent, adaptive behavior.
@@ 2 Collective Adaptive Intelligence: Conceptual Definitions
Building on the prior discussion of emergence, modularity, and adaptivity, we now formalize the concept of Collective Adaptive Intelligence (CAI). This section offers a framework for distinguishing between different levels of collective intelligence and introduces key features, adaptation, resilience, scalability, and self-assembly—that distinguish CAI from traditional multi-agent systems. These distinctions help ground future system design in both theory and practice.
@@
We first consider a conceptual definition of collective adaptive intelligent (CAI) system as shown in Figure 1. Canonical Multi-Agent Systems (MAS) are trained on single tasks and deployed solely on the training task. In traditional settings, the coordination patterns among components are predefined, which is regarded as specialist collective intelligence (SCI). However, if the MAS is pretrained on a diverse set of numerous tasks, it can potentially be deployed not only on a broad range of in-distribution tasks but also on some out-of-distribution (OOD)tasks. In such cases,it is considered as a generalist collective intelligence (GCI). CAI should possess at least the following features (as shown in Figure 1):
@@
Adaptation (CAI) vs. Zero-Shot Generalization (GCI): The difference between adaptation and zero-shot generalization is often blurred in discussions of model generalization. Yet, this distinction becomes critical in settings governed by partially observable Markov decision processes (POMDPs), where zero-shot generalization is fundamentally limited, as shown in prior work on meta-learning and general-purpose agents [9, 5]. In collective systems, where each agent only perceives a fragment of the environment and must rely on limited inter-agent communication—generalization is rarely achievable without some degree of adaptation. The need to integrate partial and dynamic information at inference time makes adaptation a more viable and often necessary strategy. Crucially, attempts to maximize zero-shot generalization can work against the flexibility required for robust performance in such decentralized, communication-bound systems. Therefore, designing truly scalable collective intelligence systems demands prioritizing adaptive mechanisms over fixed generalization.
@@ 3 A Conceptual Framework for Collective Adaptive Agents
Building on the foundational properties of collective adaptivity outlined in the previous section, we now introduce a conceptual framework for constructing Collective Adaptive Agents (CAAs)—the foundational units of Collective Adaptive Intelligence (CAI).
(a) A CAA formulation
(b) Training Process
@@
Figure 2: An illustration of the formulation of Collective Adaptive Agent (CAA) and its training process.
A well-structured conceptual framework for collective adaptive agents (CAAs) must incorporate the following core attributes to ensure robustness and scalability in distributed systems:
  • Decentralization: Agents operate independently within a decentralized system, eliminating reliance on centralized control mechanisms.
  • Self-Adaptation: Agents dynamically adjust their input-output mappings and messaging strategies based on historical data. This capability is realized through real-time adaptation of internal memory structures or learning algorithms.
  • Self-Assembly and Collective Scalability: Agents autonomously modify their communication protocols to selectively accept messages from specific peers and route outputs to relevant recipients.
As depicted in Figure 2, to fulfill the aforementioned requirements, a collective adaptive agent (CAA) must account for historical interaction data and feedback signals to dynamically adapt its role within a collective system, thereby enabling effective task generalization across diverse scenarios. The proposed CAA architecture incorporates three distinct parameter sets:
  • Static Parameters: These remain fixed during inference to maintain structural consistency.
  • Dynamic Memories: These are updated in real-time to enable self-adaptation to novel contexts or roles.
  • Communication Protocols: These alters coordination strategies among agents to align with evolving task demands.
This complex system can be optimized by drawing inspiration from meta-learning principles [9], which results in a hierarchical learning process (Figure 2(b)):
  • Outer-Loop Optimization: Searches for optimal static parameters to define the agent’s foundational capabilities and adaption functionalities.
  • Inner-Loop Adaptation: Dynamically updates dynamic memory and communication protocols during inference to ensure rapid role-switching and task-specific coordination.
Together, these design principles highlight a key insight: scalable collective intelligence emerges not from static specialization, but from dynamic, decentralized agents capable of continual adaptation in structure, behavior, and communication.
4 Limitations and Future Challenges
While collective intelligence offers a compelling paradigm for scalable embodied AI, current computational architectures are poorly suited to its decentralized, communication-heavy nature. A key challenge lies in developing Collective Adaptive Agents (CAAs) that can be trained efficiently at small scales yet generalize through self-organization at deployment. Early signs from LLMs and open-ended agents are promising, but realizing scalable, adaptive collectivity demands co-evolution of algorithms, architectures, and infrastructure. The path forward lies not in scaling monolithic models, but in designing systems where intelligence emerges through adaptive interaction.